In a jib crane, the jib and tie rod are 6 m and 5 m long respectively. The height of the crane post is 3 m. When a weight (W) 317 N is suspended from the crane and the tie rod remains horizontal.


Determine the tension in Tie rod (only magnitude, in N)?


1
Expert's answer
2021-06-07T05:49:43-0400

hr = 6m - 3m = 3m

tie rod = 5m


"\theta = tan^{-1}(\dfrac35)"

"\theta = 31°"


"\cos 31°= \dfrac{317}{T}"


"T =\dfrac{317}{\cos31°}= 369.8\ N"